OpenStudy (johnw):

Show that ∫√(2+x^2)dx = x√(2+x^2)/2 + ln(x+√(2+x^2)) + C where C is an arbitrary constant I don´t even know where to start :)

OpenStudy (518nad):

tan or sec sub

OpenStudy (518nad):

oh we are just showing, just take the derivative of the right side

OpenStudy (3mar):

May I help?

OpenStudy (518nad):

@JohnW are you there?

OpenStudy (3mar):

seems not

OpenStudy (johnw):

Yes

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Hope that helps

OpenStudy (johnw):

I´m sorry, but I don´t follow. I don´t even know what an arbitrary constant is :/

OpenStudy (johnw):

Even less knowing how to show that C is one

OpenStudy (518nad):

You are just starting derivatives?

OpenStudy (518nad):

I think they just want you to derive the right side and show this statement is true

OpenStudy (518nad):

the derivative of a constant is 0, as a derivative tells you the rate of change wrt some variable, and a constant has no change

OpenStudy (518nad):

first thing you should be aware of is that derivative is a linear operation

OpenStudy (518nad):

so that means d/dx(x√(2+x^2)/2 + ln(x+√(2+x^2)) + C) =d/dx [x√(2+x^2)/2] + d/dx [ ln(x+√(2+x^2)) ]+ d/dx[ C ]

OpenStudy (518nad):

work on deriving one term at a time, mar3 will help u i have to help someone rn

OpenStudy (johnw):

ok, thanks!

OpenStudy (3mar):

@JohnW

OpenStudy (3mar):

Like what 518nad said: Derive the right side and compare to the left side, if identical, you are correct, if not you may need help

OpenStudy (3mar):

\[\frac{ d }{ dx }(\frac{ x \sqrt{x^2+2} }{ 2 })=(\frac{ x }{ 2 })(\frac{ 1 }{ 2 })\frac{ 2x }{ \sqrt{x^2+2} }+\sqrt{x^2+2}(\frac{ 1 }{ 2 })\]

OpenStudy (3mar):

\[\frac{ d }{ dx }(\ln(x+\sqrt{x^2+2}))=\frac{ 1 }{ x+\sqrt{x^2+2} }(1+\frac{ 1 }{ 2 }\frac{ 2x }{ \sqrt{x^2+2} })\]

OpenStudy (3mar):

\[\frac{ d }{ dx }(C)=0\]

OpenStudy (3mar):

Sum all the three results, and simplify the terms

OpenStudy (johnw):

ok

OpenStudy (3mar):

Do you want to do it for you or you can do it?

OpenStudy (3mar):

@JohnW Let me know please, brother!

OpenStudy (3mar):

Hope that helps

OpenStudy (johnw):

Please help me :)

OpenStudy (johnw):

I´ll only get some of the numbers mixed up and then the entire calculation is screwed up ;)

OpenStudy (3mar):

Where are you stuck?

OpenStudy (3mar):

Brother! Where are you stuck?

OpenStudy (johnw):

I`m sorry. I have a lot going on at the same time. I think I can follow the process, but I´m not familiar with how you simplify the result.

OpenStudy (3mar):

I am busy too but when you are ready to continue, please let me know.

OpenStudy (johnw):

Thanks. now I`m online again and I`m looking for help with this question.

OpenStudy (3mar):

Sorry, I was not here!

OpenStudy (3mar):

Thank you for the medal!